# Robust Weight Imprinting: Insights from Neural Collapse and Proxy-Based Aggregation

### v3

@@ 4 Experimental Setup
We use resnet18, resnet50 (resnet), vit_b_16 (vit), and swin_b (swin) as FMs, two CNN-based and two Transformer-based architectures. All four models are pre-trained on ImageNet-1K (ILSVRC 2012) (imagenet). To generate the embeddings, we use PyTorch’s torchvision models.
@@
To find out the best imprinting strategy within our IMPRINT framework, we focus on tasks T created from the datasets MNIST (mnist), FashionMNIST (fashionmnist), and CIFAR-10 (cifar10), each containing 10 classes. We mainly focus on the three T containing all ten classes. Furthermore, we look at smaller tasks only containing classes $\{0,1,2\}$, and the two tasks containing classes $\{1,3,5,7,9\}$ resp. $\{0,2,4,6,8\}$. This random selection of $3\cdot 4=12$ tasks adds variation to our evaluations.
@@
Moreover, we construct a new out-of-distribution, non-collapsed dataset by merging all classes from four digit datasets: MNIST, MNIST-M (mnistm), SVHN (svhn), and USPS (usps). The resulting dataset, which we refer to as CombiDigits, exhibits reduced collapse due to the greater distributional diversity within each class. To ensure scale invariance in covariance-based NC measurements, all embeddings are $L^{2}$-normalized before computing $\mathcal{NC}_{1}$.
@@
Unless stated otherwise, we report the median test accuracy over three different seeds. In sections˜5.1 and 5.2, we evaluate the imprinting performance by varying the FM (4) and T (12) and report average accuracy and standard deviation (std) across these. Due to the heterogeneity of models and tasks, large std are expected. Therefore, overall method comparisons are based on ranks rather than absolute accuracies. Methods are ranked by their median accuracy, yielding $4\cdot 12=48$ potentially different ranks. We report the average rank and assess statistical significance of ranking (dis-)agreements using critical difference (CD) diagrams as explained in section˜3.3. The code used to generate these diagrams is inspired by cddiag-code.
@@ 5.1 Best Imprinting Strategy
We provide a comparison between existing memory-constrained methods used for imprinting on foundation models in fig.˜2, namely, the work by qi-imprinting; itsdone; ncmbaseline, as well as a novel, best-performing configuration (“Ours”) that results from IMPRINT. Focusing on $k=20$, we find that our method, consisting of k-means weight generation, L2 normalizations, and max aggregation, outperforms all previously studied approaches by a margin of $4\%$ on average with statistical significance. For reference, we additionally report an oracle baseline that uses cross-class feature statistics to generate its weights, representing an upper bound that is not constrained by imprinting. The results indicate that our new imprinting method significantly narrows the gap between single-proxy mean imprinting and this oracle baseline.
@@
Employing gradient-based methods – such as setting weights via the analytical least-squares initialization proposed by harun2025good, which uses data and labels from all classes jointly – can be seen as an upper bound (“Oracle” in fig.˜2) for the unsupervised weight imprinting approaches discussed here. We extend this method by combining it with k-means, forming k-least-squares, and observe in section˜A.5 that using multiple proxies per class instead of a single weight vector can still improve performance, particularly on datasets with high $\mathcal{NC}_{1}$.
@@
We restrict all the subsequent experiments to using L2 across all NORM following qi-imprinting. This combination of normalizations is chosen to specifically model cosine similarity within max aggregation.
@@ 6 Conclusion
In particular, we find that the mean weight generation (GEN) method, despite its prominence in previous work, falls short compared to k-means – even when the number of proxies $k$ is very small. Remarkably, with as little as 50 samples per class, k-means can already outperform the original imprinting method proposed by qi-imprinting, highlighting its advantage in low-data regimes.
@@
The usage of both weight and activation sparsity as shen2023reducing could change the within- and between-class variability in favor of using a higher number of proxies. Besides only using the penultimate layer embeddings for generating the classifier weights, an interesting area of study could be extracting embeddings from previous layers of the FM for this purpose. Recently, the study by marczak2025revisiting showed that adding a multi-layer perceptron projector between the penultimate and classification layers results in representations that are more transferable. Another avenue of research is to thoroughly analyze imprinting the weights of other layers as well (masked-imprinting).
@@ A.2 Datasets
MNIST-M (mnistm), SVHN (svhn), USPS (usps).
@@ A.3 Imprinting as Memory
We revisit the classical idea of Bidirectional Associative Memories (BAMs) (kohonen; anderson72; nakano; anderson77) and the associated update rule (sejnowski; dayan) for storing key-value pairs as in
@@
Notably, this form of direct memory update has seen renewed attention in modern architectures beyond standard query-key-value attention. In particular, the xLSTM model (xlstm) implements this mechanism within its mLSTM memory blocks, where the matrix memory cell is updated by gated key-value associations, closely following this classical covariance rule, indicating a broader resurgence of associative memory principles in contemporary sequence modeling.
@@ A.4 Differences between Foundation Models
Figure˜6(a), similar to fig.˜11(b), illustrates the impact of varying the number of proxies on imprinting accuracy across different foundation models (FMs). The key difference in this figure is the use of none for NORMpre instead of L2. This seemingly minor change reveals a striking contrast between CNN- and Transformer-based architectures: a distinct and consistent dip between $k=1$ and $k=d$ appears in Transformer-based models, whereas this dip is absent in fig.˜11(b), where L2 is used as NORMpre, and does not occur at all in the resnet models. We hypothesize that this difference arises from the distinct embedding distributions of CNN- and Transformer-based architectures (see, e.g., itsdone).
@@ A.5 Learned Weights (with Multiple Proxies) and Comparisons
harun2025good recently studied the initialization of classifier weights for novel categories in a Continual Learning (CL) setting using the least squares algorithm, comparing it to random initialization and class mean imprinting across various loss functions. In contrast, our work centers on imprinting-based approaches, which avoid using gradient-based optimization and cross-class statistics, operating instead on a per-class basis with immediate availability of data. Nonetheless, we include the least-squares method as a supervised oracle baseline – explicitly not an imprinting method – fine-tuned for classification accuracy.
@@
resnet18, resnet50 (resnet) 0.0001
vit_b_16 (vit) 0.1
swin_b (swin) 0.05

### v4

@@ 4 Experimental Setup
We use resnet18, resnet50 (He et al., 2016), vit_b_16 (Dosovitskiy et al., 2021), and swin_b (Liu et al., 2021) as FMs, two CNN-based and two Transformer-based architectures. All four models are pre-trained on ImageNet-1K (ILSVRC 2012) (Deng et al., 2009). To generate the embeddings, we use PyTorch’s torchvision models.
@@
To find out the best imprinting strategy within our IMPRINT framework, we focus on tasks T created from the datasets MNIST (Deng, 2012), FashionMNIST (Xiao et al., 2017), and CIFAR-10 (Krizhevsky et al., 2009), each containing 10 classes. We mainly focus on the three T containing all ten classes. Furthermore, we look at smaller tasks only containing classes $\{0,1,2\}$, and the two tasks containing classes $\{1,3,5,7,9\}$ resp. $\{0,2,4,6,8\}$. This random selection of $3\cdot 4=12$ tasks adds variation to our evaluations.
@@
Moreover, we construct a new out-of-distribution, non-collapsed dataset by merging all classes from four digit datasets: MNIST, MNIST-M (Ganin et al., 2016), SVHN (Netzer et al., 2011), and USPS (Hull, 1994). The resulting dataset, which we refer to as CombiDigits, exhibits reduced collapse due to the greater distributional diversity within each class. To ensure scale invariance in covariance-based NC measurements, all embeddings are $L^{2}$-normalized before computing $\mathcal{NC}_{1}$.
@@
Unless stated otherwise, we report the median test accuracy over three different seeds. In sections˜5.1 and 5.2, we evaluate the imprinting performance by varying the FM (4) and T (12) and report average accuracy and standard deviation (std) across these. Due to the heterogeneity of models and tasks, large std are expected. Therefore, overall method comparisons are based on ranks rather than absolute accuracies. Methods are ranked by their median accuracy, yielding $4\cdot 12=48$ potentially different ranks. We report the average rank and assess statistical significance of ranking (dis-)agreements using critical difference (CD) diagrams as explained in section˜3.3. The code used to generate these diagrams is inspired by Ismail Fawaz et al. (2019).
@@ 5.1 Best Imprinting Strategy
We provide a comparison between existing memory-constrained methods used for imprinting on foundation models in fig.˜2, namely, the work by Qi et al. (2018); Hosoda et al. (2024); Janson et al. (2022), as well as a novel, best-performing configuration (“Ours”) that results from IMPRINT. Focusing on $k=20$, we find that our method, consisting of k-means weight generation, L2 normalizations, and max aggregation, outperforms all previously studied approaches by a margin of $4\%$ on average with statistical significance. For reference, we additionally report an oracle baseline that uses cross-class feature statistics to generate its weights, representing an upper bound that is not constrained by imprinting. The results indicate that our new imprinting method significantly narrows the gap between single-proxy mean imprinting and this oracle baseline.
@@
Employing gradient-based methods – such as setting weights via the analytical least-squares initialization proposed by Harun and Kanan (2025), which uses data and labels from all classes jointly – can be seen as an upper bound (“Oracle” in fig.˜2) for the unsupervised weight imprinting approaches discussed here. We extend this method by combining it with k-means, forming k-least-squares, and observe in section˜A.5 that using multiple proxies per class instead of a single weight vector can still improve performance, particularly on datasets with high $\mathcal{NC}_{1}$.
@@
We restrict all the subsequent experiments to using L2 across all NORM following Qi et al. (2018). This combination of normalizations is chosen to specifically model cosine similarity within max aggregation.
@@ 6 Conclusion
In particular, we find that the mean weight generation (GEN) method, despite its prominence in previous work, falls short compared to k-means – even when the number of proxies $k$ is very small. Remarkably, with as little as 50 samples per class, k-means can already outperform the original imprinting method proposed by Qi et al. (2018), highlighting its advantage in low-data regimes.
@@
The usage of both weight and activation sparsity as Shen et al. (2023) could change the within- and between-class variability in favor of using a higher number of proxies. Besides only using the penultimate layer embeddings for generating the classifier weights, an interesting area of study could be extracting embeddings from previous layers of the FM for this purpose. Recently, the study by Marczak et al. (2025) showed that adding a multi-layer perceptron projector between the penultimate and classification layers results in representations that are more transferable. Another avenue of research is to thoroughly analyze imprinting the weights of other layers as well (Siam et al., 2019).
@@ A.2 Datasets
MNIST-M (Ganin et al., 2016), SVHN (Netzer et al., 2011), USPS (Hull, 1994).
@@ A.3 Imprinting as Memory
We revisit the classical idea of Bidirectional Associative Memories (BAMs) (Kohonen, 2009; Anderson, 1972; Nakano, 2007; Anderson et al., 1977) and the associated update rule (Sejnowski, 1977; Dayan and Willshaw, 1991) for storing key-value pairs as in
@@
Notably, this form of direct memory update has seen renewed attention in modern architectures beyond standard query-key-value attention. In particular, the xLSTM model (Beck et al., 2024) implements this mechanism within its mLSTM memory blocks, where the matrix memory cell is updated by gated key-value associations, closely following this classical covariance rule, indicating a broader resurgence of associative memory principles in contemporary sequence modeling.
@@ A.4 Differences between Foundation Models
Figure˜6(a), similar to fig.˜11(b), illustrates the impact of varying the number of proxies on imprinting accuracy across different foundation models (FMs). The key difference in this figure is the use of none for NORMpre instead of L2. This seemingly minor change reveals a striking contrast between CNN- and Transformer-based architectures: a distinct and consistent dip between $k=1$ and $k=d$ appears in Transformer-based models, whereas this dip is absent in fig.˜11(b), where L2 is used as NORMpre, and does not occur at all in the resnet models. We hypothesize that this difference arises from the distinct embedding distributions of CNN- and Transformer-based architectures (see, e.g., Hosoda et al. (2024, Figure S2)).
@@ A.5 Learned Weights (with Multiple Proxies) and Comparisons
Harun and Kanan (2025) recently studied the initialization of classifier weights for novel categories in a Continual Learning (CL) setting using the least squares algorithm, comparing it to random initialization and class mean imprinting across various loss functions. In contrast, our work centers on imprinting-based approaches, which avoid using gradient-based optimization and cross-class statistics, operating instead on a per-class basis with immediate availability of data. Nonetheless, we include the least-squares method as a supervised oracle baseline – explicitly not an imprinting method – fine-tuned for classification accuracy.
@@
resnet18, resnet50 (He et al., 2016) 0.0001
vit_b_16 (Dosovitskiy et al., 2021) 0.1
swin_b (Liu et al., 2021) 0.05